I'm trying to work the following problem by using the elimintion method. Here is the problem:
4a+3b=16
-4a+b=16
4b=32
b=8
4a+3(8)=16 At this point I'm lost and can not figure out (a).
Request your help.

Answer by jim_thompson5910(35256)

4a+3b=16
-4a+b=16
----------
0a+4b=32

4b = 32

b = 32/4

b = 8

4a + 3b = 16

4a + 3(8) = 16

4a + 24 = 16

4a = 16-24

4a = -8

a = -8/4

a = -2

So the solutions are a = -2 and b = 8